# DotResize: Reducing LLM Width via Discrete Optimal Transport-based Neuron Merging

### v1

@@ 4.1 Models
We focus our evaluation on three different families of LLMs: Llama 3.1 at 8B and 70B parameters (grattafiori2024llama, ), Mistral at 7B (v0.3) and 12B (NeMo) parameters (yang2024qwen2, ), and Phi-4 at 12B parameters (abdin2024phi, ). These models all reflect large Transformer models pre-trained on primarily English text data. We obtain all models using their huggingface transformers library implementations (wolf-etal-2020-transformers, ). All models with the exception of Llama-3.1-70B can be compressed and evaluated on a single 32GB V100 GPU, whereas we use 8 V100s for Llama-3.1-70B experiments.
@@ 4.2 Evaluation
For language modeling capabilities, we evaluate our models by computing perplexity on the English Wikitext-2 dataset (merity2017pointer, ). For zero-shot capabilities, we test the models on 5 evaluation sets sourced from lm-evaluation-harness (eval-harness, ): ARC-Challenge and ARC-Easy (clark2018think, ), HellaSwag (zellers2019hellaswag, ), PIQA (bisk2020piqa, ), and Winogrande sakaguchi2021winogrande . These five tasks reflect diverse and commonly used English tasks for evaluating language models. Following Ashkboos et al. ashkboos2024slicegpt , in all language modeling experiments, we use a small subset ($\sim$262K tokens) of the Wikitext-2 training data to compute our transportation maps. For all zero-shot experiments, we use the same number of tokens, but from the Stanford Alpaca dataset for instruction following alpaca .
4.3 Baseline Comparisons
We compare two versions of our method to two corresponding baselines. The first baseline is a simpler magnitude pruning baseline where at each layer, we prune the neurons with the lowest average $\ell_{2}$ norm across the calibration data. These neuron indices can differ across each layer by projecting residual connections as well as in our method, as we found this to be superior to whole-model single-index pruning, as proposed in LLM-Pruner as “channel-pruning” ma2023llm . When comparing our method to this baseline, Magnitude Pruning, we select these same neurons as the target support of DotResize, to get a direct comparison of pruning versus OT-based merging in this setting. This means that to select the neurons that comprise the $d_{\text{new}}$ support, we keep the neurons with highest average $\ell_{2}$ norm.
The second baseline is SliceGPT with default parameters, which first applies PCA to reproject features, and then prunes neurons with least total variance ashkboos2024slicegpt . We directly compare our method to SliceGPT by reprojecting activations by their PCA basis before applying DotResize  to reduce neuron width. We differ from their method by forgoing the slicing matrix and instead using our transport map in its place. In this version, which we denote as PCA+DotResize, we use $\ell_{1}$ norm to select neurons. Unless stated otherwise, we use $\lambda=0.1$ for the Sinkhorn regularization value.
5 Results
5.1 Language Modeling
Table 1: Perplexity performance of models on Wikitext-2 under different width reduction strategies and at different levels of width reduction (sparsity). Perplexity results that reflect > 0.1 PPL reduction are bolded. In comparing Magnitude Pruning to DotResize  without pre-projection, we see a stark improvement in perplexity, indicating that using optimal transport to reduce neuron width can be superior to pruning neurons.
Sparsity Method Llama-3.1 Mistral Phi-4
@@ 5.1 Language Modeling
Table 1 presents language modeling results for the uncompressed models, existing compression baselines, and both variants of our method. Compared to Magnitude Prune, DotResize  achieves significantly lower perplexity by preserving and leveraging signal from neurons that would otherwise be discarded. While Magnitude Prune only retains the neurons with the highest $\ell_{2}$ activation norms, DotResize  uses these high-norm neurons as a support set and redistributes the full neuron width onto them based on similarity. Although this redistribution can sometimes reduce the influence of neurons that would have been retained by Magnitude Prune, the advantage of preserving signal from the pruned neurons more than compensates for this effect.
In comparing the DotResize  results and PCA+DotResize  results, we see that firstly decorrelating the activations according to PCA improves performance. By firstly transforming the activations so that their dimensions are uncorrelated, first applying PCA can help DotResize  more effectively identify meaningful similarities based on distance in our transportation map finding step.
@@
In comparing the PCA+DotResize  results, we see that for the Llama-3.1 8B model, our method reduces the perplexity by almost 3 points at 30% width reduction, and over 1 point for 20% width reduction. For the other models, while Mistral-7B does not improve from using PCA+DotResize, other models either maintain similar performance or improve with PCA+DotResize. Given some improved performance with the difference in objectives to reduce neuron width, we are able to demonstrate an interesting limitation of SliceGPT.
@@
Using PCA to reduce neuron width in SliceGPT is the $\ell_{2}$-optimal strategy for linear dimensionality reduction, as it solves the $\ell_{2}$ reconstruction problem between original and down-projected activations. However, by outperforming SliceGPT in a comparable setting, we can demonstrate that minimizing $\ell_{2}$ activation distance does not always correlate with improved downstream performance. This indicates that relying on $\ell_{2}$ activation distance as a proxy for model performance, which is common in compression applications dong2017learning ; frantar2022optimal , may miss other aspects of model behavior.
5.2 Zero-Shot performance on downstream applications
Given that PCA+DotResize  outperforms DotResize  without pre-projection in language modeling, we evaluate only SliceGPT and PCA+DotResize  on zero-shot tasks. Accuracy results across all 5 tasks and an average accuracy are found in Table 2. As seen in the table, several model/sparsity combinations are improved using PCA+DotResize , including Mistral-Nemo 12B, which sees an over 11% average accuracy improvement across tasks at 20% sparsity. Additionally, using PCA+DotResize  improves results for Phi-4, which is remarkably robust to both our compression method and SliceGPT. At 20% sparsity, this model retains 98% of its original average zero-shot performance, and even sees higher accuracy on the ARC-Easy and Challenge sets after compression.
We also observe model-specific differences in both language modeling and zero-shot performance, with some models performing on par with or slightly worse than SliceGPT. These variations potentially stem from training differences across state-of-the-art LLMs, which can affect their compatibility with merging-based compression. In practice, we recommend evaluating both approaches and, where possible, lightly tuning our method, which is relatively inexpensive due to its transform-in-place nature.
@@ 5.2 Zero-Shot performance on downstream applications
Table 2: Zero shot performance (Accuracy) of models on downstream tasks under different width removal settings. Bolded results reflect the higher result on model and sparsity comparisons.
@@ 5.3 Analysis
Figure 3: Sparsity vs performance tradeoff across different levels of compression on Llama-3.1-70B. Sparsity (%) indicates how much width is removed from weight matrices. While 10% sparsity does not improve compute cost due to residual parameterization, after 20%, real world compute cost reduction can be observed.
As our method is intended to reduce memory use and improve latency, we measure the total compute used in inference: We finding the minimum number of 32GB V100 GPUs our compressed model can run on without GPU-offloading, and measure the latency in ms/token during inference on sequences of 128 tokens ashkboos2024slicegpt . This gives us the total compute cost in GPUms/token. We compress Llama-3.1-70B at different levels of width reduction, and evaluate both perplexity and inference cost, and report this trade-off in in Figure 3. After 20% sparsity, compute cost is reduced from the original model, dictating the region of practical use of width reduction methods.
@@
As stated, we select $\lambda=0.1$ to serve as the regularization parameter found in Equation 1. To understand the effect of different values of $\lambda$ on our method, we test three of our models on 5 different $\lambda$ values, ranging from 0.01 to 10. As seen in Figure 4, the amount of regularization used does not change downstream language modeling performance to a large extent, suggesting that there are several acceptable weighting strategies for combining neurons successfully. However, different values of $\lambda$ can result in some perplexity differences in the Llama-3.1-8B model as seen at $\lambda=0.01$. Given training differences, different values of $\lambda$ may be optimal for different models, but the values presented here reflect an acceptable range.
@@
Much like previous work that examines the behavior of neurons to dictate neuron alignment or pruning singh2020model ; ainsworthgit ; ashkboos2024slicegpt , our method uses a small amount of relevant calibration data ( 262K tokens) to determine the behavior of individual neurons. We evaluate the sensitivity of our method to the amount of calibration data used by testing the use of only 65K tokens, 131K tokens, 262K tokens, and 524K tokens to compute transportation maps. Results are shown in Figure 5.
Figure 4: Performance of models on Wikitext-2 at 20% and 30% width reduction with different amounts of entropic regularization applied to the Sinkhorn objective. Besides some notable performance loss at $\lambda=0.01$ for Llama-3.1-8B at both 20% and 30% sparsity, our method is not sensitive to the amount of entropic regularization present for these models.
Figure 5: Performance of models compressed with DotResize  on Wikitext-2 at 20% sparsity, with varying amounts of exemplar tokens used to compute transport maps in our method. After approximately 130K tokens, the returns on using more calibration data appear diminishing.
While DotResize  is not notably sensitive to the amount of calibration data used, we recommend using at least $2^{17}$ or 131K tokens for computing transport maps. This robustness is a strength, as computing the cost matrices $C$ from Equation 1 can become computationally expensive with large datasets. Additionally, this robustness suggests that a small amount of exemplar data is sufficient to capture meaningful neuron behavior, as reflected in the generalization ability of our compressed models.
@@ 6 Conclusion and Future Work
In this work, we introduce DotResize, a novel model compression method that frames neuron width reduction as an optimal transport problem in order to combine a layer’s neurons based on their activation signatures, specifically avoiding the pruning of any particular neuron. In taking this approach, we are able to demonstrate improved performance across several LLMs compared to pruning-based alternatives, demonstrating the importance of incorporating the full signal of a layer and the potential insufficiencies of pruning. Some models are hugely receptive to the addition of a merging-based objective in width reduction, with Phi-4 maintaining almost $98\%$ of zero-shot performance on downstream tasks while improving on some individual tasks at 20% sparsity, and Mistral-Nemo achieving an over 10% gain in average accuracy across tasks over SliceGPT. We also demonstrate an extension of computational invariance in Transformers by demonstrating that invertible matrices can be applied to Transformer weights via a QR-decomposition step, and without changing the model function. While we focus on model compression as our application, our work also provides a basis for exploring discrete optimal transport as a general-purpose parameter transformation method applicable beyond compression.
@@ Appendix A Simple explanation of QR Decomposition invariance
Appendix B Limitations
In discussing the limitations of this work, there are both limitations of the method proposed as well as limitations of the validation of this method. Focusing firstly on the method proposed, we acknowledge that while our method does lead to real-world compression and latency improvements, the compression attained is not comparable to methods like low-bit quantization. However, in practice, this method is orthogonal to methods like quantization and should be combined. Additionally, we acknowledge that despite being a training-free method, our method is not data-free, as we require the use of some amount of data to observe neuron behavior when computing alignments and transportation maps.
Focusing on the validation of this method, there are limitations in the scope of models tested in this work. To make a focused contribution, we limit the scope to LLMs trained primarily on English text data. However, our method can be easily generalized to other architectures, and can be applied to pre-trained Transformer models with multimodal and/or multilingual capabilities. We leave this exploration for future work.
Appendix C Broader Impacts

### v2

@@ 4.1 Models
We focus our evaluation on three different families of LLMs: Llama 3.1 at 8B and 70B parameters (Grattafiori et al., 2024), Mistral at 7B (v0.3) and 12B (NeMo) parameters (Yang et al., 2024), and Phi-4 at 14B parameters (Abdin et al., 2024). These models all reflect large Transformer models pre-trained on primarily English text data. We obtain all models using their HuggingFace transformers library implementations (Wolf et al., 2020). All models with the exception of Llama-3.1-70B can be compressed and evaluated on a single 32GB V100 GPU, whereas we use 8 V100s for Llama-3.1-70B experiments.
@@ 4.2 Evaluation
For language modeling capabilities, we evaluate our models by computing perplexity on the English Wikitext-2 dataset (Merity et al., 2017). For zero-shot general capabilities, we test the models on 5 evaluation sets sourced from lm-evaluation-harness (Gao et al., 2024a): ARC-Challenge and ARC-Easy (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), and Winogrande (Sakaguchi et al., 2021). These five tasks reflect diverse and commonly used English tasks for evaluating language models. Following Ashkboos et al. (2024a), in all language modeling experiments, we use a small subset (128 samples) of the Wikitext-2 training data to compute our transportation maps. Samples are length 2048. For all zero-shot experiments, we use the same number of tokens, but from the Stanford Alpaca dataset for instruction following (Taori et al., 2023).
Table 1: Perplexity performance of models on Wikitext-2 under different width reduction strategies and at different levels of width reduction. Perplexity results that reflect >> 0.1 PPL reduction are bolded. In comparing Magnitude Pruning to DotResize  without pre-projection, we see a stark improvement in perplexity, indicating that using optimal transport to reduce neuron width can be superior to solely pruning neurons.
Width Reduction Method Llama-3.1 Mistral Phi-4
@@
4.3 Pruning Comparisons
We apply DotResize  atop of two pruning methods. The first method is a simple magnitude-based pruner where at each Transformer sublayer, neurons with the lowest average $\ell_{2}$ norm on calibration data are pruned. Neuron indices can differ across each layer by projecting residual connections too, as detailed in section 3.3, as we found this to be superior pruning the same indices throughout the entire model, as proposed in LLM-Pruner as “channel-pruning” (Ma et al., 2023). When adding our method to this baseline, Magnitude Pruning, we select the non-pruned neurons as the target support of DotResize, to get a direct comparison of pruning versus OT-based merging in this setting. This means that to select the neurons that comprise the $d_{\text{new}}$ support, we keep the neurons with highest average $\ell_{2}$ norm.
The second pruning method is SliceGPT, which first applies PCA to reproject features, and then prunes neurons with least total variance (Ashkboos et al., 2024a). This method reflects a popular and competitive approach within width pruning. We directly compare our method to SliceGPT by reprojecting activations by their PCA basis before applying DotResize  to reduce neuron width, and call this method PCA+DotResize. Neurons are selected by $l_{1}$ norm. We differ from their method by forgoing the slicing matrix and instead using our transport map in its place. We use Sinkhorn regularization $\lambda=0.1$ for both settings.
5 Results
5.1 Language Modeling
@@
Table 1 presents language modeling results for the uncompressed models, existing pruning techniques, and our method applied atop the pruning methods. In applying DotResize atop Magnitude Prune, DotResize substantially lowers model perplexity by preserving and leveraging signal from neurons that would otherwise be discarded. While Magnitude Prune only retains the neurons with the highest $\ell_{2}$ activation norms, DotResize  uses these high-norm neurons as a support set and redistributes the full neuron width onto them based on similarity. Although this redistribution can sometimes reduce the influence of neurons that would have been retained by Magnitude Prune, the advantage of preserving signal from the pruned neurons more than compensates for this effect, as seen by the perplexity reduction.
@@
In comparing the PCA+DotResize  results with SliceGPT, we see that for the Llama-3.1 8B model, our method reduces the perplexity by almost 3 points at 30% width reduction, and over 1 point for 20% width reduction. For the other models, while Mistral-7B does not improve from using PCA+DotResize, other models either maintain similar performance or improve with PCA+DotResize. Given that DotResize can improve performance with the difference in objectives to reduce neuron width, we are able to demonstrate an interesting limitation of SliceGPT.
@@
Using PCA to reduce neuron width in SliceGPT is the $\ell_{2}$-optimal strategy for linear dimensionality reduction, as it solves the $\ell_{2}$ reconstruction problem between original and down-projected activations. However, by outperforming SliceGPT in a comparable setting, we can demonstrate that minimizing $\ell_{2}$ activation distance does not always correlate with improved downstream performance. This indicates that relying on $\ell_{2}$ activation distance as a proxy for model performance, which is common in compression applications (Dong et al., 2017; Frantar & Alistarh, 2022), may miss other aspects of model behavior.
@@
Table 2: Zero shot performance (% Accuracy) of models on downstream tasks under different width removal settings. Bolded results reflect the higher result on model and sparsity comparisons.
@@
5.2 Zero-Shot performance on downstream applications
Given that PCA+DotResize outperforms Magnitude Pruning-based DotResize in language modeling, we evaluate only SliceGPT and PCA+DotResize on zero-shot tasks. Accuracy results across all 5 tasks and an average accuracy are found in Table 2. As seen in the table, several model/sparsity combinations are improved using PCA+DotResize, including Mistral-Nemo 12B, which sees a substantial 11% average accuracy improvement across tasks at 20% sparsity. Additionally, using PCA+DotResize improves results for Phi-4 at 20%, which is remarkably robust to both our compression method and SliceGPT. At 20% sparsity, this model retains 98% of its original average zero-shot performance, and even sees higher accuracy on the ARC-Easy and Challenge sets after compression versus in the base model.
We also observe model-specific differences in both language modeling and zero-shot performance, with certain models being very receptive to this signal recombination approach, and others benefiting only slightly or not at all. These variations potentially stem from training differences across state-of-the-art LLMs, which can affect their compatibility with merging-based compression. To investigate this further, we plot the effective rank of activations (at 99% variance) for Llama-3.1-8B, enhanced by DotResize, and Mistral-7B, which follows SliceGPT-like trends. These results, captured during compression, are detailed in Appendix C. We find that Mistral activations have a much lower effective rank than Llama activations in earlier layers, suggesting that Mistral activations contain most of its signal within its first few neurons after PCA reprojection. Despite this concentration, we find our method remains sensitive to varying-width reduction strategies across layers, mirroring findings in Ashkboos et al. (2024a). In practice, we recommend evaluating both approaches and, where possible, lightly tuning our method, which is relatively inexpensive due to its transform-in-place nature.
@@ 5.3 Analysis
Figure 3: Sparsity vs performance tradeoff across different levels of compression on Llama-3.1-70B. Sparsity (%) indicates how much width is removed from weight matrices. After 20%, real world compute cost reduction can be observed.
Figure 4: Performance on Wikitext-2 at 20% and 30% width reduction with different entropic regularization $\lambda$. Our method is not sensitive to the amount of regularization for these models.
Figure 5: Performance at 20% sparsity with varying exemplar tokens. After approximately 130K tokens, the returns on using more calibration data appear diminishing.
As our method is intended to reduce memory use and improve latency, we measure the total compute used in inference: we find the minimum number of 32GB V100 GPUs the compressed Llama-3.1-70B model can run on without GPU-offloading, and measure the latency in ms/token during inference on sequences of 128 tokens (Ashkboos et al., 2024a). This gives us the total compute cost in GPUms/token. We compress at different levels of width reduction, and evaluate both perplexity and inference cost, and report this trade-off in in Figure 5. After 20% width reduction, the overall latency begins to improve; the small region of increased cost is incidental to the cost involved in reprojecting the residual stream in both SliceGPT and our method.
@@
We select $\lambda=0.1$ as the regularization parameter in Equation 1. To understand the effect of different values of $\lambda$ on our method, we evaluate three of our models on 5 different $\lambda$ values, ranging from 0.01 to 10. As seen in Figure 5, the amount of regularization does not significantly impact downstream language modeling performance, suggesting that multiple weighting strategies can successfully recombine neurons. However, minor variations in $\lambda$ can result in some perplexity differences in the Llama-3.1-8B model as seen at $\lambda=0.01$. While optimal $\lambda$ values may vary due to training differences across models, the results presented here indicate a broad range of effective settings.
@@
Similar to prior work that examines the behavior of neurons for alignment or pruning (Singh & Jaggi, 2020; Ainsworth et al., 2023; Ashkboos et al., 2024a), our method uses a small amount of relevant calibration data (128 samples) to determine the behavior of individual neurons. We evaluate the sensitivity of our method to the amount of calibration data used by testing the use of 32 samples, 64 samples, 128 tokens, and 256 samples to compute transportation maps. Results are shown in Figure 5.
While DotResize  is not notably sensitive to the amount of calibration data used, we recommend using at least 64 samples for computing transport maps. This robustness is a strength, as computing the cost matrices $C$ from Equation 1 can become computationally expensive with large datasets. Additionally, this robustness suggests that a small amount of exemplar data is sufficient to capture meaningful neuron behavior, as reflected in the generalization ability of our compressed models.
@@ 6 Conclusion and Future Work
In this work, we introduce DotResize, a novel model compression method that frames neuron width reduction as an optimal transport problem in order to re-project a layer’s neurons, specifically avoiding the pruning of any particular neuron. In taking this approach, we are able to demonstrate improved performance across several LLMs compared to pruning-based alternatives, demonstrating the importance of incorporating the full signal of a layer and the potential insufficiencies of pruning. We find some models to be hugely receptive to the addition of a merging-based objective in width reduction, with Phi-4 maintaining almost $98\%$ of zero-shot performance on downstream tasks while improving on some individual tasks at 20% sparsity, and Mistral-Nemo achieving an over 10% gain in average accuracy across tasks over SliceGPT. We also demonstrate a technical extension of computational invariance in Transformers by demonstrating that invertible matrices can be applied to Transformer weights via a QR-decomposition step, and without changing the model function. Several directions merit future exploration. The model-specific variations we observe suggest that intrinsic model properties may influence OT-based compressibility, understanding these factors could enable more adaptive compression strategies. More broadly, discrete optimal transport offers a principled framework for parameter transformation that may extend beyond compression.
Impact Statement
@@ Appendix A Simple explanation of QR Decomposition invariance
Appendix B QR Decomposition invariance for rectangular maps
We will prove a theorem related to the RMSNorm, and then discuss what the result dictates in terms of invariance to RMSNorm in the rectangular transport map $T$ case.
Theorem B.1 (QR invariance for rectangular maps).
Let $T\in\mathbb{R}^{d_{\text{orig}}\times d_{\text{new}}}$ be QR-decomposed into $T=QR$, where semi-orthogonal $Q\in\mathbb{R}^{d_{\text{orig}}\times d_{\text{new}}}$ has orthonormal columns ($Q^{T}Q=I_{d_{\text{new}}}$) and $R\in\mathbb{R}^{d_{\text{new}}\times d_{\text{new}}}$ is upper triangular and invertible. Then,
$\text{RMSNorm}(xQ)RT^{\dagger}=\frac{xQQ^{T}}{\|xQQ^{T}\|}.$ (11)
Proof.
We first note that $Q$ is semi-orthogonal, therefore
$\|xQ\|$ $=\sqrt{xQQ^{T}x^{T}}=\sqrt{xQ(Q^{T}Q)Q^{T}x^{T}}=\|xQQ^{T}\|.$ (12)
Next, using the QR decomposition of $T$, the Moore–Penrose pseudoinverse is
$T^{\dagger}$ $=(T^{T}T)^{-1}T^{T}$
$=(R^{T}Q^{T}QR)^{-1}R^{T}Q^{T}$
$=(R^{T}R)^{-1}R^{T}Q^{T}$
$=R^{-1}Q^{T}.$
Substituting into the expression of interest gives
$\text{RMSNorm}(xQ)RT^{\dagger}$ $=\frac{xQ}{\|xQ\|}R(R^{-1}Q^{T})$ (13)
$=\frac{xQQ^{T}}{\|xQ\|}$ (14)
$=\frac{xQQ^{T}}{\|xQQ^{T}\|},$ (15)
which completes the proof. ∎
Given this result, we see that the RMSNorm is invariant to our constructed maps up to orthogonal projection onto the column space of $Q$.
Appendix C Layer-wise analysis of effective rank
For a data matrix of $n$ items of dimension $d_{\text{orig}}$, we compute effective rank as the following, assuming we have ordered eigenvalues $\{\sigma_{i}\}_{i\in[1,d_{\text{orig}}]}$, and threhold $0<\tau<1$:
$k=\arg\min_{k^{\prime}}\frac{\sum_{i=1}^{k^{\prime}}\sigma_{i}}{\sum_{i=1}^{d_{\text{orig}}}\sigma_{i}}>\tau$
In other words, we find the minimum index $k$ such that the cumulative sum of the first $k$ eigenvalues exceeds ratio $\tau$. In this analysis, we fix $\tau=0.99$. We compute the effective rank for two models, Llama-3.1-8B and Mistral-7B across all 32 layers, at both the output of the attention sublayer and the output of the MLP sublayer. Both models have true model dimension of 4096. We report our results in Figure 6, and find that Mistral has significantly reduced effective rank for many early layers as compared to Llama, suggesting that Llama has more disperse signal across its PCA-reprojected neurons.
Figure 6: Effective rank at 99% variance of Llama-3.1-8B and Mistral-7B activations at each layer during PCA+DotResize application for 20% width reduction. We report values for both the output of attention and the MLP, after the residual connection. We find that Mistral reaches a much smaller effective rank in earlier layers, suggesting that it contains less distributed signal across neurons after PCA reprojection.